OpenStudy (markchernioglo):

Write the slope-intercept equation of the line that has the same slope as y-3=1/2(x+3) and contains the point (8,4). Explain how you wrote the equation.

OpenStudy (markchernioglo):

I know I use y=mx+b

OpenStudy (daniellelovee):

replace the x with 8 and the y with 4

OpenStudy (markchernioglo):

Oh so 4-3=1/2(8+3) ?

OpenStudy (daniellelovee):

yup

OpenStudy (markchernioglo):

http://prntscr.com/cpx1ue

OpenStudy (markchernioglo):

Thanks, but look at the answer they give..

OpenStudy (daniellelovee):

you already have the slope which is 1/2

OpenStudy (markchernioglo):

They want it in some wierd way lol.

OpenStudy (daniellelovee):

ok give me what they show you

OpenStudy (markchernioglo):

http://prntscr.com/cpx1ue

OpenStudy (markchernioglo):

Even though I have the answer I need the explanation because I want to pass this upcoming test. :)

OpenStudy (daniellelovee):

oh yeah I was just going to say that the pink letter where the answer xD

OpenStudy (daniellelovee):

ok so what don't you understand?

OpenStudy (markchernioglo):

Well how they got the 9/2

OpenStudy (markchernioglo):

I tried to understand but nope I can't lol

OpenStudy (daniellelovee):

ok I'm going to do it the way I would do it and maybe that is easier for you to understand

OpenStudy (markchernioglo):

Alright. The outcome is still y=1/2x?

OpenStudy (daniellelovee):



OpenStudy (daniellelovee):

so far so good?

OpenStudy (markchernioglo):

Oh yes sorry.

OpenStudy (daniellelovee):

is ok :)

OpenStudy (daniellelovee):

did you understand?

OpenStudy (markchernioglo):

yea makes sense the way you did it!

OpenStudy (daniellelovee):

alright great! any other questions?

OpenStudy (markchernioglo):

I think they want me to use the formula y=mx+b, but thank you very much! I got no other questions atm. :)

OpenStudy (daniellelovee):

yes they want you to change from point slope intercept formula into slope intercept formula